the diffence between the length and width of one side of a rectangular cereal box is 9.8cm if the width is 21.8cm what is the length in cm?
how many centimeters long is the perimeter of the rectangle?

Answer by HyperBrain(694)

Let x=the width
y=the length
(1)y-x=9.8
(2)x=21.8
Substituting (2) to(1),
y-21.8=9.8
y=31.6
Therefore, the length is 31.6 cm